OpenStudy (anonymous):

Okay this is the next one, I'm thinking it looks easy, so I could try it on my own first then you could give me your feedback! :) Question: Let f(x) = −4x + 7 and g(x) = 10x − 6. Find f(g(x)). Answer Choices: A) -40x + 24 B) -40x + 31 C) -40x + 64 D) -40x + 70

OpenStudy (anonymous):

@Maddy1251

OpenStudy (maddy1251):

Okay, go ahead and solve it, I'll solve it on my own on this side, and we will go from there. :-)

OpenStudy (anonymous):

Uhh.. I thought it'd be easy. But I noticed there isn't a substitution for the variables lol. So, when I combined the polynomials, I didn't know where to go from there. Subtract 6 from both sides maybe? xD

OpenStudy (anonymous):

This is what I got with the combining by the way: 10x - 6 - 4x + 7.

OpenStudy (maddy1251):

I'll get you started to help out, and see if you can get it :-) This is just like the one before, expect we don't need to solve for f(g(x)), we just need to find what it equals. So we have:\ Let f(x) = −4x + 7 and g(x) = 10x − 6. Find f(g(x)) Okay, lets start this. we need to find f(g(x)). What does g(x)=? 10x-6. Great. Now lets look at f(x), where it = -4x+7. Now, 10x-6 is our 'x' for f(x), since we are solving for f(g(x)), or if you really want to think of it, f(x). Because g(x)=x for f(x). Now, what will you do with 10x-6, when looking at -4x+7? You got a little confused but on the right track.

OpenStudy (anonymous):

I understood everything you've said so far. So g(x) is actually x for f(x). And I'm guessing that means we'd have to take 10x-6 and plug it into the -4x+7 but since it equals x we'd place it between the -4 and the +7?

OpenStudy (maddy1251):

Yep! 10x-6 would go in for 'x', looking like this: -4(10x-6)+7.

OpenStudy (maddy1251):

f(g(x))=f(10x-6)=-4x+7=-4(10x-6)+7

OpenStudy (anonymous):

Distribute then right? The -4 would multiply into the 10x and the -6?

OpenStudy (maddy1251):

Right. What will you get?

OpenStudy (anonymous):

-40x + 24 + 7?

OpenStudy (maddy1251):

Yes! Now, add 24 and 7.

OpenStudy (anonymous):

The answer would become B: -40x + 31 !! Well this was easy lol. The other one I just didn't know where the equations went. Lol. Thank you for the help Maddy. :)

OpenStudy (anonymous):

Would you be up for another one?

OpenStudy (maddy1251):

Excellent job! -claps- I am SO happy for you! See, you got this :) I am up for more if you are!

OpenStudy (anonymous):

Of course, I'll close this question and tag you in the next one! :)

OpenStudy (maddy1251):

Great :-)